The Edwards family and the Powell family each used their sprinklers last summer. The water output rate for the Edwards family's sprinkler was
25L
per hour. The water output rate for the Powell family's sprinkler was
20L
per hour. The families used their sprinklers for a combined total of
55
hours, resulting in a total water output of
1300L
. How long was each sprinkler used?

Answer by addingup(3677)

Let's call 25 x and 20 y:
x+y= 55 Subtract x from both sides:
y= 55-x We will use this value for y in the next equation
------------
25x+20y= 1300 Now substitute for the value of y:
25x+20(55-x)= 1300
25x+1100-20x= 1300 Subtract 1100 on both sides and add/subtract x on left:
5x= 200 Divide both sides by 5:
x= 40 This is how many hours the Edwards sprinklers ran at 25l per hour.
55-40= 15 hours at 20l per hour (Powell family)
Proof:
40*25+15*20= 1300 We have the correct answer